OpenStudy (anonymous):

Which fraction falls between three-tenths and three-fifths? five-tenths eight-tenths four-fifths two-tenths

OpenStudy (anonymous):

five tenths because

OpenStudy (anonymous):

three-tenths =3/10 and three-fifths=3/5=6/10 (upon multiplying by two) hence the only number between this is 5/10 ie five-tenths

OpenStudy (anonymous):

any more doubts?

OpenStudy (anonymous):

\[\frac{3}{10}<\frac{4}{10}<\frac{5}{10}<\frac{6}{10} \]